OpenStudy (anonymous):

Need some help, this is most confusing to me then anything!! MEDALS After a discount of 20%, the price of a pair of shoes, including tax, is $66.91. The following equation can be used to find x, the price of the pair of shoes without the discount: x - 0.2x + 0.06(x - 0.2x) = 66.91 What is the price of the pair of shoes without the discount?

OpenStudy (anonymous):

$78.90

OpenStudy (anonymous):

can you explain why? just so i know how to do it in the future?

OpenStudy (anonymous):

x-0.2x = 66.91/1.06 = 63.1226 x = 63.1226/0.8 = 78.90

OpenStudy (anonymous):

medals and fan

OpenStudy (anonymous):

pweese

OpenStudy (anonymous):

best response

OpenStudy (anonymous):

coo

OpenStudy (anonymous):

become fan ?